3-19 Moms Motel Corporation (MM) plans to issue bonds to raise $175 million that it needs to support future operations. MMs investment banker will charge 2.5 percent of the total amount issued to help MMM raise the funds. In addition, MM will incur other costs associated with the issue that total $500,000. The market value of each bond at issue time will be $1,000. How many bonds must MM sell to net $175 million after flotation costs? Assume that fractions of bonds cannot be issued. Show how much of the issue will consist of flotation costs and how much MM will receive after flotation costs are paid. 3-20 United Uninsured Underwriters (U3) needs $192 million to stay in business. If it issues new common stock to raise the funds, the flotation costs will be 8 percent. The new issue will also require U3 to pay $280,000 in fees to its lawyers, printing costs, and other costs associated with the issue. U3 can issue stock at $25 per share. How many shares of common stock must U3 issue so that it has $192 million after flotation costs? Show how much of the total dollar amount of the issue will be flotation costs and how much U3 will receive after the flotation costs are paid.
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

Home nert Page Layout Formulas Data Review View dd-Ins Cut Σ AutoSum ー E ゴWrap Text ta copy. в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Sort &Find & 2 ClearFe Select Edting Format Painter Formatting, as Table w styles. ▼ ㆆ ▼ Clipboard Alignment Number Cells DG DH DI DJ DK DL DM DN DO DP DR DS DT DU 3-19 AMOUNT REQUIRED OTHER ISSUE COST THIS IS THE AMOUNT TO BE RAISED BEFORE 2.5% COST- 175 MILLION 0.5 MILLION 175.5 MILLION 4 5 6 SO 100-2.5-97.5% THEN 100%= 175.5 180 9 SO AMOUNT TO BE RAISED = 180 MILLION 10 NO OF BONDS TO BE ISSUED = 180 MILLION/1000 = 180000 12 13 14 15 16 17 18 19 TOTAL FLOTATION COST- OUT OF WHICH 2.5% OF INVESTMENT BANKER- OTHER ISSUE COST 5 MILLION 4.5 MILLION 0.5 MILLION 180-175 = (180*2.5%) MM WILL RECEIVE NET 175 MILLION YIELD SPOT Sheet2 AFN BANKING NOTE BIDDING, UNDERWRITING EUAC AW PM LIFE LP I MRP. INFLATION YIELD WARRANT fund bond -CLEAN INVOICE PRICE LI 福 130% 31-01-2019Home nert Page Layout Formulas Data Review View dd-Ins Cut E AutoSum Calibri ー E ゴWrap Text General aCopy в 1 프 . Ej-., Δ. : rーー 逻锂函Merge & Center. $, % , 弼,8 Conditional Format eCell Insert Delete Format Paste Sort &Find & Format Painter Formatting as Table Styles2 Clear Clipboard Font Alignment Number Styles Cells Edting DP33 DG DH DI DJ DK DL DM DN DO DP DQ DR DS DT 18 19 20 21 3--20 AMOUNT REQUIRED OTHER ISSUE COST THIS IS THE AMOUNT TO BE RAISED BEFORE 8% COST- 192 MILLION 0.28 MILLION 192.28 MILLION 23 24 25 26 27 28 29 30 31 32 SO 100-8 = 92% THEN 100% 192.28 209 SO AMOUNT TO BE RAISED = 209 MILLION NO OF BONDS TO BE ISSUED = 209 MILLION/25 = 8360000 TOTAL FLOTATION COST = OUT OF WHICH 8% IS FLOTATION COST- OTHER ISSUE COST 209 192 17 16.72 MILLION 0.28 MILLION (209*8%) 34 35 36 COMPANY WILL RECEIVE NET 192 MILLION YIELD SPOT Sheet2 AFN BANKING NOTE BIDDING, UNDERWRITING EUAC AW PM LIFE LP I MRP. INFLATION YIELD WARRANT fund bond -CLEAN INVOICE PRICE LI 06:10 31-01-2019